mathslover (mathslover):

The function \(f(x)=x^4−10x^3+40x^2−80x+64\) has four complex roots, one of which is \(2−2i\). What is the sum of all real and imaginary coefficients of these roots?

OpenStudy (anonymous):

if one i 2-2i then one more will be 2+2i

mathslover (mathslover):

Should I proceed like this? : \((2-2i)\times z = f(x) \) ?

mathslover (mathslover):

And then find the value of z by dividing? ... Ok! @shubham.bagrecha

mathslover (mathslover):

So I can do : \((2-2i)(2+2i)(z) = f(x)\)

OpenStudy (anonymous):

other two will also be conjugates

mathslover (mathslover):

How can we say?

OpenStudy (anonymous):

you said that all four are complex and complex roots occur in conjugate forms

mathslover (mathslover):

Ok.

mathslover (mathslover):

I am getting another factor as : \(\cfrac{1}{8}\) . (real factor) .

mathslover (mathslover):

hmn I am wrong somewhere .

OpenStudy (anonymous):

let the four roots be : 2-2i ,2+2i ,a+bi ,a-bi from product of roots taken one at a time , a^2 + b^2 = 8

mathslover (mathslover):

Yep.

OpenStudy (anonymous):

4+2a=10 a=3 (sum of roots)

mathslover (mathslover):

Ok so I have : 3 + bi and 3 -bi yet as two another factors.

mathslover (mathslover):

I can calculate \(3^2 + b^2 = 8\) \(b^2 =1\) \(b = \pm 1\)

OpenStudy (anonymous):

b^2 = -1

mathslover (mathslover):

Oh yes , sorry : \(b = i\)

OpenStudy (anonymous):

so something is wrong , b can't be imaginery

mathslover (mathslover):

a^2 + b^2 = 8 There is wrong in ^ above.

mathslover (mathslover):

Well how did you get 8 there?

OpenStudy (anonymous):

no it is correct

OpenStudy (anonymous):

which means our roots are : 2+2i,2-2i,2,4

mathslover (mathslover):

That is we don't have more complex roots than 2+2i and 2-2i

OpenStudy (anonymous):

yes

mathslover (mathslover):

So 10 is what I haev

OpenStudy (anonymous):

and the other roots are satisfying f(x)

mathslover (mathslover):

^ have.

mathslover (mathslover):

Got it. Thanks @shubham.bagrecha . Good work :)

OpenStudy (anonymous):

thanks